Sensor Networks for Structural Monitoring: Status, Plans, Problems

Ananth Grama

Goal

Designing sensing infrastructure for real-time, high fidelity physical measurement retrieval

Test infrastructure: three story 30′ × 50′ structure

• Physical phenomenon to be measured:

– Linear displacement

– Multi-axial acceleration

Linear Displacement Sensing

– Structural strain • Four commonly-used options:

– Optical tracking

– Time of flight and triangulation sensors

– Magnetic, inductive and capacitive sensors

– Strain gauges and resistive sensors • Optical tracking

– Mechanism ∗ Markers on the surface

∗ Cameras capture mark displacement

∗ Image processing calibrated to measure distance – Pros

∗ Non-contact sensing

∗ Simple and safe installation

– Cons ∗ Expensive video capture equipment

∗ Difficult calibration

∗ High processing requirements

∗ High data transfer rates • Time of flight and triangulation sensors

– Mechanism ∗ Sound or light (laser or IR) waves incident on the surface ∗ Measures round trip time or phase shift of the wave to

measure distance

– Pros

∗ Laser sensors have high accuracy (up to few microns) and wide measurement ranges (up to few feet)

∗ Ultrasonic sensors are safe, most laser sensors in the market are eye-safe

∗ Most products usually ship with RS-232 I/O interface – Cons

∗ Ultrasonic sensors do not have high accuracy and depend on ambient conditions

∗ Laser sensors are very expensive (≈ $2, 000)

∗ Requires 12V DC power supply • Magnetic, inductive and capacitive sensors

– Mechanism

∗ Magnetic and inductive sensors use the relative movement of a soft iron core and a magnetic field to measure displacements

∗ Capacitive sensors measure capacitance changes as distance or overlapping area between conductive plates change

– Pros ∗ These sensors are very cheap

– Cons

∗ Capacitive sensors and some magnetic sensors have a very short range and span of distances

∗ Most magnetic and inductive sensors with higher ranges involve the movement of concentric rods (one attached to the moving surface and other to the fixed support)

∗ Sensors are affected by ambient conditions like temperature and humidity • Strain gauges and resistive (potentiometer) sensors

– Mechanism

∗ A shaft in contact with the surface moves, resulting in electrical resistance change which is translated to distance

∗ Another design has an extensible chord and a mountable base, whereby pulling the chord results in change of resistance of a potentiometer in the base

– Pros ∗ These sensors are very cheap

∗ The chord version can extend up to thousands of inches

∗ Essentially the resolution is infinite, however resistance variability may cause inaccuracies

∗ Chorded version also allows velocity measurement

– Cons ∗ Involves moving parts implying low life due to wear

∗ Need external circuitry and calibration to translate voltage differences to distance

∗ Requires 25V supply; variation in input voltage may also cause inaccuracies • Use of a few laser sensors for high precision measurements

and several chorded potentiometers appear to be an attractive option.

• Battery packs can be used to supply power, however the life expectancy of such power packs needs to be evaluated.

• Chorded potentiometers can also be used to evaluate structural strain

Multi-axial acceleration

• 2-axis and 3-axis accelerometers are available

• Most options have are MEMs devices with small size and low power requirements

• Most common sensors have +/- 2g range and with +/- 2 mg sensitivity, and are available for only tens of dollars

• Higher end accelerometers are available for a couple of hundred dollars

• Example: MTS420CA is a data acquisition board with 2-axis accelerometer, developed by UC Berkeley and Intel and is compatible with MICA2

Network Plan

• For our current test infrastructure both wired and wireless options are possible

• Wireless option allows easier deployment but lower data rates

• For wireless 433 MHz radio communication equipment is available, this allows a maximum of 42 packets per second (12.364 Kbps)

• Using a high number of sensing devices over a large geographical area favors an ad-hoc network of nodes

• Fewer devices in a small area can all directly report to a single master gateway (forming a star topology)

Network Plan: continued..

• MICA2 mote is a good possibility because it also allows programmability and on board processing capability

• For a high density deployment an overlay of gateways with one interface to the mote wireless network and the other to RS-232, Ethernet, or 802.11b can be deployed

• Data streams from the gateways can be redirected via commodity networks to the data sinks possibly over campus networks, WANs, or even the Internet

• Bluetooth communication devices are also an option

Implementation Plan

• In the next few weeks, we must have complete deployment in preparation for tests.

• Set up displacement sensors first – displacement at periphery is adequate at this point.

• Perhaps, start from optical sensing and see if we can use laser rangefinders? (Eyesafe lasers are difficult to scale to larger distances).

• Accelerometers are easy to deploy. These can be done at short notice.

Research and Implementation Issues

• Data rate

• Errors

• Calibrations